According to the index monday.com china produces the largest amount of ammonia with 45,000 528000 metric tonne the profitability of the ammonia production is around 90% of the total ammonia production costs which is about 10,000 perton refer to the balanced chemical equation N2 + 3H2 = 2nh3 how much is the production cost of this amount of ammonia


1
Expert's answer
2021-11-08T08:14:37-0500

m(H2)=m(NH3)×3Mr(H2)/2Mr(NH3)=45520×3×2/(2×17)=8,033 thousand metric ton;


m(N2)=m(NH3) - m(H2)=45,520-8,033=37,487 thousand metric ton;